3/5 plus 1/10 = ?
5/9 - 1/3= ?
1/3 plus 2/5 = ?

Answer by ewatrrr(24785)

3/5 plus 1/10 = 6/10 + 1/10 = 7/10
5/9 - 1/3= 5/9 - 3/9 = 2/9
1/3 plus 2/5 = 5/15 + 6/15 = 11/15